# Competency-Aware Planning for Probabilistically Safe Navigation Under Perception Uncertainty

### v4

@@ I INTRODUCTION
With the prevalence of DNN models in many real-world applications and concern about their failure modes, there has been extensive research on quantifying uncertainty in these models [6] and detecting inputs that are outside of their training distribution [7]. This has enabled the development of uncertainty-aware navigation schemes that explicitly consider the existence of uncertainty arising from DNN-based perception models. However, existing navigation schemes generally rely on uncertainty estimation methods that do not perform well for data outside of the model’s training distribution or on methods that provide binary decisions about whether an input is anomalous based on the training set. These methods do not allow for the design of probabilistically safe control schemes that maintain performance under multiple aspects of predictive model uncertainty.
@@ III PROBLEM FORMULATION
In this work, we consider a UGV navigating a simulated lunar environment. During training, the perception model learns to distinguish between different regions in the environment (e.g., bumpy terrain, smooth terrain, inside crater, edge of crater, etc.) to navigate effectively. During evaluation, it encounters astronauts and human-made obstacles that were not seen during training. These unfamiliar obstacles are outlined in white in Figure 5 and are considered OOD because they were not present in the training set. The vehicle is expected to recognize regions in the environment that are unfamiliar to the perception model and continue to navigate effectively in the face of predictive uncertainty. To enable effective navigation, we develop a probabilistic competency estimation method for the perception model (Section IV), then use these competency estimates to develop competency-aware planning and control schemes (Section V). Additional details and parameters are provided in Appendices -A and -B.
@@ IV-A Estimating Overall Model Competency
$\hat{p}_{D|C}$ $=P\bigl{(}\{\mathcal{L}_{c}>\ell(\bm{X})-(\mu_{c}+z\sigma_{c})\}\bigr{)}$
$=1-P\bigl{(}\{\mathcal{L}_{c}\leq\ell(\bm{X})-(\mu_{c}+z\sigma_{c})\}\bigr{)}$
$=1-F_{\mathcal{L}_{c}\sim\mathcal{N}(\mu_{c},\sigma_{c})}\bigl{(}\ell(\bm{X})-\mu_{c}-z\sigma_{c}\bigr{)}$
$=1-F_{Z\sim\mathcal{N}(0,1)}\left(\frac{\ell(\bm{X})-2\mu_{c}-z\sigma_{c}}{\sigma_{c}}\right)$
$=:1-\phi\left(\frac{\ell(\bm{X})-2\mu_{c}}{\sigma_{c}}-z\right).$
@@
$\hat{\rho}(\bm{X}):=\hat{p}_{\hat{c}}\sum_{c\in\mathcal{C}}\hat{p}_{c}\left(1-\phi\left(\frac{\ell(\bm{X})-2\mu_{c}}{\sigma_{c}}-z\right)\right).$
We refer to $\hat{\rho}(\bm{X})$ as the overall competency score for image $\bm{X}$. If the model is 100% competent on the image, the probability that its prediction is correct is one. If it is entirely incompetent, the probability its prediction is correct is zero. The competency score is thus between zero and one, with higher scores corresponding to higher levels of competency.
@@ IV-B Analyzing Overall Model Competency Scores
We compare our overall model competency estimation method to twelve existing methods for UQ and OOD detection. We find that, while the capability of our method to distinguish between correctly classified and misclassified samples is comparable to existing methods, our scoring method is much better at distinguishing between ID and OOD samples. In addition, we believe our scoring method is more intuitive than others, producing scores that are always between zero and one, as opposed to existing methods with no preset range. See Appendix -C for a more thorough comparison of our competency score to existing methods.
@@ IV-D Analyzing Regional Model Competency Maps
We compare our regional competency estimation method to seven existing methods for anomaly detection and localization. We find that our method outperforms existing methods when tasked with distinguishing between familiar and unfamiliar pixels, but its performance is not significantly better than that of PaDiM [45]. However, compared to PaDiM, which produces anomaly scores in a variable range, our method generates probabilistic scores between 0 and 1, which are more interpretable and straightforward to incorporate into a probabilistic planning framework. See Appendix -D for a more thorough comparison of our regional model competency maps to existing methods.
@@ V COMPETENCY-AWARE NAVIGATION
We use the overall model competency estimates (§IV-A) and regional competency estimates (§IV-C) to design navigation schemes with varying levels of competency awareness. We develop a model of the vehicle dynamics (§V-A), use this model to design a competency-aware path planner (§V-B), and develop a path-tracking controller (§V-C).
V-A Model of Vehicle Dynamics
We model our vehicle as a non-linear discrete time system whose state, $\vec{x}\in\mathbb{R}^{5}$, can be described in terms of its X-position ($x$), Y-position ($y$), heading ($\theta$), linear velocity ($v$), and turn rate ($\omega$). The input, $\vec{u}\in\mathbb{R}^{2}$ to the system is the desired linear velocity or throttle ($t$) and desired turn rate or steering command ($s$). The state and input are expressed as
$\vec{x}=\begin{bmatrix}x&y&\theta&v&\omega\end{bmatrix}^{T}\hskip 5.69054pt\text{and}\hskip 5.69054pt\vec{u}=\begin{bmatrix}t&s\end{bmatrix}^{T}\hskip 5.69054pt\text{respectively.}$.
The vehicle is then described by a discrete time-varying model that is very similar to the Dubin’s car model:
$\vec{x}_{k+1}=\bm{A_{k}}\vec{x}_{k}+\bm{B}\vec{u}_{k},\ \text{where}$
$\bm{A_{k}}=\begin{bmatrix}1&0&0&\Delta t\cos{\tilde{\theta}_{k}}&0\\ 0&1&0&\Delta t\sin{\tilde{\theta}_{k}}&0\\ 0&0&1&0&\Delta t\\ 0&0&0&1-\alpha&0\\ 0&0&0&0&1-\beta\end{bmatrix}\hskip 14.22636pt\bm{B}=\begin{bmatrix}0&0\\ 0&0\\ 0&0\\ \alpha&0\\ 0&\beta\\ \end{bmatrix}.$
In the above equation, $\Delta t$ is the time step, and the parameters $\alpha$ and $\beta$ are estimated from data. To generate a linear approximation for the vehicle dynamics, we use an estimate of the vehicle’s heading, $\tilde{\theta}_{k}$, at future time steps.
@@ V-A Model of Vehicle Dynamics
V-B Competency-Aware Path Planning
@@
V-C Path-Tracking Controller
To follow the paths generated by the competency-aware path planner, we develop a reference-tracking linear-quadratic regulator (LQR) using the linearized model of vehicle dynamics from Section V-A. The reference state and input are obtained from the competency-aware path planner.
V-D Evaluation of Navigation Performance
@@
We conduct ten trials in each of the five navigation scenarios shown in Figure 5. We evaluate navigation performance based on: (1) Success rate: percentage of trials in which the vehicle reaches its goal within the allotted time. (2) Timeout rate: percentage of trials in which the vehicle fails to reach the goal in the allotted time. (3) Collision rate: percentage of trials in which the vehicle collides at least once. (4) Navigation time: average time required to navigate to the goal (in seconds). (5) Path length: average distance traveled by the vehicle to its goal (in meters). The results of these experiments are summarized in Table I.
@@
To see a visual comparison of the baseline controller, the controller using only overall competency, the turning-based controller utilizing both overall and regional competency, and the trajectory-based controller utilizing both types of competency information, please view our demo video. [^2] A visual comparison of navigation performance across different planners is available on YouTube: https://youtu.be/mUB2MDQZObU.
@@ VII LIMITATIONS & FUTURE WORK
While the integration of competency-awareness into the navigation framework greatly reduces the number of collisions with unfamiliar obstacles, collisions do still occur for a couple reasons. The field-of-view of the vehicle’s front-facing camera is quite limited, and the vehicle currently has no visual memory, so it will collide with obstacles that it cannot see. It would be useful for the vehicle to maintain some memory of competency estimates for nearby regions that are no longer visible. The competency estimators are also imperfect, and errors in competency estimates carry through the perception, planning, and control pipeline, which can result in collisions. It would be interesting to explore how these errors propagate through the navigation pipeline. As a final note, we are currently employing a relatively simple graph-based image segmentation algorithm, whose limitations can affect the performance of our regional competency estimator. Future work could explore domain-specific segmentation methods to enhance performance.
There are also various other avenues for future work. First, there is currently an implicit assumption that spatial features of an image, corresponding to unfamiliar obstacles, cause the reduction in perception model competency. It would be useful to develop a more general framework for understanding model competency in cases where predictive uncertainty is not due to some particular region in the image/environment. Second, the competency estimates are currently formulated for classification problems. One could generalize these formulations to be used for other tasks. In addition, this work focused on integrating competency-awareness into a local planning and control framework. It would be interesting to develop a global planner with competency-awareness. Finally, the current response to perception model competency is to avoid regions associated with low competency. To enable continual learning within a dynamic or unfamiliar environment, it would be useful to intentionally and cautiously explore unfamiliar regions, rather than simply avoid them. It would be interesting to consider methods for safe exploration and decision frameworks to trade off between avoidance and exploration.

### v5

@@ I INTRODUCTION
With the prevalence of DNN models in many real-world applications and concern about their failure modes, there has been extensive research on quantifying uncertainty in these models [6] and detecting inputs that are outside of their training distribution [7]. This has enabled the development of uncertainty-aware navigation schemes that explicitly consider the existence of uncertainty arising from DNN-based perception models. However, existing navigation schemes generally rely on uncertainty estimation methods that do not perform well for data outside of the model’s training distribution [8] or on methods that provide binary decisions about whether an input is anomalous based on the training set [9]. These methods do not allow for the design of probabilistically safe control schemes that maintain performance under multiple aspects of predictive model uncertainty.
@@ III PROBLEM FORMULATION
In this work, we consider a UGV navigating a simulated lunar environment, in which robust visual navigation is very advantageous. During training, the perception model learns to distinguish between different regions in the environment (e.g., bumpy terrain, smooth terrain, inside crater, edge of crater, etc.) to navigate effectively. During evaluation, it encounters astronauts and human-made obstacles that were not seen during training. These unfamiliar obstacles are outlined in white in Figure 5 and are considered OOD because they were not present in the training set. The vehicle is expected to recognize regions in the environment that are unfamiliar to the perception model and continue to navigate effectively in the face of predictive uncertainty. To enable effective navigation, we develop a probabilistic competency estimation method for the perception model (Section IV), then use these competency estimates to develop competency-aware planning and control schemes (Section V). Additional details and parameters are provided in Appendices -A and -B.
@@ IV-A Estimating Overall Model Competency
$\hat{p}_{D|C}$ $=P\bigl{(}\{\mathcal{L}_{c}>\ell(\bm{X})-(\mu_{c}+z\sigma_{c})\}\bigr{)}$ (6)
$=1-P\bigl{(}\{\mathcal{L}_{c}\leq\ell(\bm{X})-(\mu_{c}+z\sigma_{c})\}\bigr{)}$ (7)
$=1-F_{\mathcal{L}_{c}\sim\mathcal{N}(\mu_{c},\sigma_{c})}\bigl{(}\ell(\bm{X})-\mu_{c}-z\sigma_{c}\bigr{)}$ (8)
$=1-F_{Z\sim\mathcal{N}(0,1)}\left(\frac{\ell(\bm{X})-2\mu_{c}-z\sigma_{c}}{\sigma_{c}}\right)$ (9)
$=:1-\phi\left(\frac{\ell(\bm{X})-2\mu_{c}}{\sigma_{c}}-z\right).$ (10)
@@
$\hat{\rho}(\bm{X}):=\hat{p}_{\hat{c}}\sum_{c\in\mathcal{C}}\hat{p}_{c}\left(1-\phi\left(\frac{\ell(\bm{X})-2\mu_{c}}{\sigma_{c}}-z\right)\right).$ (11)
We refer to $\hat{\rho}(\bm{X})$ as the overall competency score for image $\bm{X}$. If the model is 100% competent on the image, the probability that its prediction is correct is one. If it is entirely incompetent, the probability its prediction is correct is zero. The competency score is thus between zero and one, with higher scores corresponding to higher levels of competency [^2] Additional details about this competency score and how it is calibrated are available on arXiv: https://arxiv.org/abs/2411.16715..
@@ IV-B Analyzing Overall Model Competency Scores
We compare our competency estimator to twelve existing methods for UQ and OOD detection. The capability of our method to distinguish between correctly classified and misclassified samples is comparable to existing methods, but our scoring method is much better at distinguishing between ID and OOD samples. Comparing our method to competing ones, the Kolmogorov–Smirnov (KS) distance between ID and OOD distributions is two times greater or more. When seeking to identify OOD samples from a set of ID samples, our method achieves an area under the receiver operating characteristic curve (AUROC) score of nearly one, and the false positive rate (FPR) is close to zero in both cases, which is an order of magnitude smaller than existing methods. In addition, we believe our scoring method is more intuitive than others, producing scores that are always between zero and one, as opposed to existing methods with no preset range. See Appendix -C for a more thorough comparison of our competency score to existing methods.
@@ IV-D Analyzing Regional Model Competency Maps
We compare our regional competency estimation method to seven existing methods for anomaly detection and localization. We find that our method outperforms existing methods when tasked with distinguishing between familiar and unfamiliar pixels, achieving the lowest FPRs, the highest KS distance between familiar and unfamiliar pixels in OOD images, the second highest KS distance between pixels in ID images and unfamiliar pixels in OOD images, and the second highest AUROC values. Although its performance is not significantly better than that of PaDiM [45], PaDiM produces anomaly scores in a variable range, while our method generates probabilistic scores between 0 and 1, which are more interpretable and straightforward to incorporate into a planning framework. See Appendix -D for a more thorough comparison of our regional model competency maps to existing methods.
@@ V COMPETENCY-AWARE NAVIGATION
We use the overall model competency estimates (§IV-A) and regional competency estimates (§IV-C) to design navigation schemes with varying levels of competency awareness (§V-A) and evaluate them across five scenarios (§V-B).
@@
V-A Competency-Aware Path Planning
@@
To follow the paths generated by the competency-aware path planner, we develop a reference-tracking linear-quadratic regulator (LQR) using a linearized model of vehicle dynamics. See Appendix -A for more details on our vehicle model, competency-aware path planner, and path-tracking controller.
V-B Evaluation of Navigation Performance
@@
We conduct ten trials in each of the five navigation scenarios shown in Figure 5. All of these scenarios are implemented in a simulated lunar environment built on Gazebo, and all evaluations are run in real time. We evaluate navigation performance based on: (1) Success rate: percentage of trials in which the vehicle reaches its goal within the allotted time. (2) Timeout rate: percentage of trials in which the vehicle fails to reach the goal in the allotted time. (3) Collision rate: percentage of trials in which the vehicle collides at least once. (4) Navigation time: average time required to navigate to the goal (in seconds). (5) Path length: average distance traveled by the vehicle to its goal (in meters). The results of these experiments are summarized in Table I.
@@
A more nuanced evaluation of navigation schemes across different scenarios is provided in Appendix -E. To see a visual comparison of the baseline controller, the controller using only overall competency, the turning-based controller utilizing both overall and regional competency, and the trajectory-based controller utilizing both types of competency information, please view our demo video.333 A visual comparison of navigation performance across different planners is available on YouTube: https://youtu.be/mUB2MDQZObU.
@@ VII LIMITATIONS & FUTURE WORK
While the integration of competency-awareness into the navigation framework greatly reduces the number of collisions with unfamiliar obstacles, collisions do still occur for a couple reasons. The field-of-view of the vehicle’s front-facing camera is quite limited, and the vehicle currently has no visual memory, so it will collide with obstacles that it cannot see. It would be useful for the vehicle to maintain some memory of competency estimates for nearby regions that are no longer visible. The competency estimators are also imperfect, and errors in competency estimates carry through the perception, planning, and control pipeline, which can result in collisions. It would be interesting to explore how these errors propagate through the navigation pipeline. As a final note, we are currently employing a relatively simple graph-based image segmentation algorithm, whose limitations can affect the performance of our regional competency estimator. Future work could explore domain-specific segmentation methods or emerging foundation models for image segmentation to enhance performance.
There are also various other avenues for future work. First, there is currently an implicit assumption that spatial features of an image, corresponding to unfamiliar obstacles, cause the reduction in perception model competency. It would be useful to develop a more general framework for understanding model competency in cases where predictive uncertainty is not due to some particular region in the image/environment. Second, the competency estimates are currently formulated for classification problems. One could generalize these formulations to be used for other tasks. In addition, this work focused on integrating competency-awareness into a local planning and control framework. It would be interesting to develop a global planner with competency-awareness. Additionally, the current response to perception model competency is to avoid regions associated with low competency. To enable continual learning within a dynamic or unfamiliar environment, it would be useful to intentionally and cautiously explore unfamiliar regions, rather than simply avoid them. It would be interesting to consider methods for safe exploration and decision frameworks to trade off between avoidance and exploration. As a final note, while robust visual navigation is critical in lunar environments, robustness to perception uncertainty is important in other domains as well. One might consider the application of this work to other areas of field robotics.